OpenStudy (anonymous):

Approximate the instantaneous rate of change for the function. f(x)=square root of x at x=4

OpenStudy (anonymous):

What sort of approximation?

OpenStudy (anonymous):

All I know this is given for instantaneous rate of change lim b->a f(b)-f(a)/b-a

OpenStudy (anonymous):

That isn't an approximation though. That's the exact formula.

OpenStudy (anonymous):

\[ \lim_{x\to 4}\frac{\sqrt x -\sqrt 4}{x-4} \]

OpenStudy (anonymous):

Multiply top and bottom by conjugate of numerator. Can you do that?

OpenStudy (phi):

do you know how to take the derivative, or must you use lim b->a f(b)-f(a)/b-a ?

OpenStudy (anonymous):

I am not sure how, but it looks I need to use lim b->a f(b)-f(a)/b-a

OpenStudy (phi):

if you use the limit formula, it looks like this. Let a= 4, and b = x+h= 4+h \[\lim_{h \rightarrow 0}\frac{ \sqrt{4+h }-\sqrt{4 }}{ (4+h)-4 }\]

OpenStudy (anonymous):